# Semantic Integrity Constraints: Declarative Guardrails for AI-Augmented Data Processing Systems

### v1

@@ 8. Conclusions
SICs are a unified abstraction that extend traditional database integrity constraints to enhance the reliability of AI-augmented DPSs. They provide a declarative, relational interface to easily specify common classes of constraint while maintaining the flexibility to support more sophisticated use cases. Crucially, this declarative interface enables the DPS to apply a wide range of highly optimized implementations. By adopting this principled approach, SICs open new, grounded research avenues towards the goal of leveraging AI-augmented DPSs for critical workloads where reliability is paramount.
@@ Acknowledgements.
- AI (2024) Guardrails AI. 2024. Detect Jailbreak. https://hub.guardrailsai.com/validator/guardrails/detect_jailbreak
@@
- DeepSeek-AI (2025) DeepSeek-AI. 2025. DeepSeek-R1: Incentivizing Reasoning Capability in LLMs via Reinforcement Learning. arXiv:2501.12948 [cs.CL] https://arxiv.org/abs/2501.12948
@@
- Goyal and Durrett (2020) Tanya Goyal and Greg Durrett. 2020. Evaluating Factuality in Generation with Dependency-level Entailment. In Findings of the Association for Computational Linguistics: EMNLP 2020, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, Online, 3592–3603. https://doi.org/10.18653/v1/2020.findings-emnlp.322
@@
- IBM (2023) IBM. 2023. What is dark data? Retrieved 2025-02-27 from https://www.ibm.com/think/topics/dark-data
@@
- Kryscinski et al. (2020) Wojciech Kryscinski, Bryan McCann, Caiming Xiong, and Richard Socher. 2020. Evaluating the Factual Consistency of Abstractive Text Summarization. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), Bonnie Webber, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, Online, 9332–9346. https://doi.org/10.18653/v1/2020.emnlp-main.750
- Laban et al. (2022) Philippe Laban, Tobias Schnabel, Paul N. Bennett, and Marti A. Hearst. 2022. SummaC: Re-Visiting NLI-based Models for Inconsistency Detection in Summarization. Transactions of the Association for Computational Linguistics 10 (2022), 163–177. https://doi.org/10.1162/tacl_a_00453
- Levy et al. (2024) Mosh Levy, Alon Jacoby, and Yoav Goldberg. 2024. Same Task, More Tokens: the Impact of Input Length on the Reasoning Performance of Large Language Models. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Lun-Wei Ku, Andre Martins, and Vivek Srikumar (Eds.). Association for Computational Linguistics, Bangkok, Thailand, 15339–15353. https://doi.org/10.18653/v1/2024.acl-long.818
@@
- Liu et al. (2025) Chunwei Liu, Matthew Russo, Michael Cafarella, Lei Cao, Peter Baille Chen, Zui Chen, Michael Franklin, Tim Kraska, Samuel Madden, Rana Shahout, and Gerardo Vitagliano. 2025. Palimpzest: Optimizing AI-Powered Analytics with Declarative Query Processing. In CIDR.
@@
- Llama (2024) Meta Llama. 2024. Prompt Guard. https://www.llama.com/docs/model-cards-and-prompt-formats/prompt-guard/
@@
- Madden et al. (2024) Samuel Madden, Michael Cafarella, Michael Franklin, and Tim Kraska. 2024. Databases Unbound: Querying All of the World’s Bytes with AI. Proc. VLDB Endow. 17, 12 (Aug. 2024), 4546–4554. https://doi.org/10.14778/3685800.3685916
- Microsoft (2024) Microsoft. 2024. Prompt Shields. https://learn.microsoft.com/en-us/azure/ai-services/content-safety/concepts/jailbreak-detection
- OpenAI (2024a) OpenAI. 2024a. Learning to Reason with LLMs. https://openai.com/index/learning-to-reason-with-llms/
- OpenAI (2024b) OpenAI. 2024b. Structured Outputs. https://platform.openai.com/docs/guides/structured-outputs/
@@
- Park et al. (2024) Kanghee Park, Jiayu Wang, Taylor Berg-Kirkpatrick, Nadia Polikarpova, and Loris D'Antoni. 2024. Grammar-Aligned Decoding. In Advances in Neural Information Processing Systems, A. Globerson, L. Mackey, D. Belgrave, A. Fan, U. Paquet, J. Tomczak, and C. Zhang (Eds.), Vol. 37. Curran Associates, Inc., 24547–24568. https://proceedings.neurips.cc/paper_files/paper/2024/file/2bdc2267c3d7d01523e2e17ac0a754f3-Paper-Conference.pdf
@@
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, brian ichter, Fei Xia, Ed Chi, Quoc V Le, and Denny Zhou. 2022. Chain-of-Thought Prompting Elicits Reasoning in Large Language Models. In Advances in Neural Information Processing Systems, S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, and A. Oh (Eds.), Vol. 35. Curran Associates, Inc., 24824–24837. https://proceedings.neurips.cc/paper_files/paper/2022/file/9d5609613524ecf4f15af0f7b31abca4-Paper-Conference.pdf
@@
- Zha et al. (2023) Yuheng Zha, Yichi Yang, Ruichen Li, and Zhiting Hu. 2023. AlignScore: Evaluating Factual Consistency with A Unified Alignment Function. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, Toronto, Canada, 11328–11348. https://doi.org/10.18653/v1/2023.acl-long.634

### v2

@@ 8. Conclusions
SICs offer a unified, declarative foundation for improving the reliability of AI-augmented data processing. By generalizing traditional integrity constraints to the semantic behaviors of LLM-augmented operators, SICs bridge the gap between declarative query processing and the uncertainty of generative models. Their integration into the planning and execution layers of modern DPSs enables not only reliable enforcement but also cost-aware optimization, system-wide coordination, and organizational reuse. Through this framework, we outline a principled path toward building auditable, trustworthy AI-augmented systems capable of supporting mission-critical workflows. SICs are not only a guardrail mechanism, but also a catalyst for a new class of optimizations, abstractions, and research directions at the intersection of databases and LLMs.
@@ Acknowledgements.
- Aiken et al. (1992) Alexander Aiken, Jennifer Widom, and Joseph M Hellerstein. 1992. Behavior of database production rules: Termination, confluence, and observable determinism. In Proceedings of the 1992 ACM SIGMOD international conference on Management of data. 59–68.
@@
- Anthropic (2025) Anthropic. 2025. Claude’s extended thinking. https://www.anthropic.com/news/visible-extended-thinking
@@
- Chern et al. (2023) I-Chun Chern, Steffi Chern, Shiqi Chen, Weizhe Yuan, Kehua Feng, Chunting Zhou, Junxian He, Graham Neubig, and Pengfei Liu. 2023. FacTool: Factuality Detection in Generative AI–A Tool Augmented Framework for Multi-Task and Multi-Domain Scenarios. arXiv:2307.13528 [cs.CL] https://arxiv.org/abs/2307.13528
@@
- Dai et al. (2024) Hanjun Dai, Bethany Yixin Wang, Xingchen Wan, Bo Dai, Sherry Yang, Azade Nova, Pengcheng Yin, Phitchaya Mangpo Phothilimthana, Charles Sutton, and Dale Schuurmans. 2024. UQE: A Query Engine for Unstructured Databases. In Advances in Neural Information Processing Systems, A. Globerson, L. Mackey, D. Belgrave, A. Fan, U. Paquet, J. Tomczak, and C. Zhang (Eds.), Vol. 37. Curran Associates, Inc., 29807–29838. https://proceedings.neurips.cc/paper_files/paper/2024/file/34b3a40ec9752c1ae48fe85fef8fe8dc-Paper-Conference.pdf
- Dayal et al. (1988) Umeshwar Dayal, Barbara Blaustein, Alex Buchmann, Upen Chakravarthy, Meichun Hsu, R Ledin, Dennis McCarthy, Arnon Rosenthal, Sunil Sarin, Michael J. Carey, et al. 1988. The HiPAC project: Combining active databases and timing constraints. ACM Sigmod Record 17, 1 (1988), 51–70.
@@
- Franklin et al. (2011) Michael J. Franklin, Donald Kossmann, Tim Kraska, Sukriti Ramesh, and Reynold Xin. 2011. CrowdDB: answering queries with crowdsourcing. In Proceedings of the 2011 ACM SIGMOD International Conference on Management of Data (Athens, Greece) (SIGMOD ’11). Association for Computing Machinery, New York, NY, USA, 61–72. https://doi.org/10.1145/1989323.1989331
- Google (2025) Google. 2025. Gemini 2.5 brings enhanced reasoning to enterprise use cases. https://cloud.google.com/blog/products/ai-machine-learning/gemini-2-5-pro-flash-on-vertex-ai
@@
- Guo et al. (2017) Chuan Guo, Geoff Pleiss, Yu Sun, and Kilian Q. Weinberger. 2017. On calibration of modern neural networks. In Proceedings of the 34th International Conference on Machine Learning - Volume 70 (Sydney, NSW, Australia) (ICML’17). JMLR.org, 1321–1330.
- Hanson (1996) Eric N. Hanson. 1996. The design and implementation of the Ariel active database rule system. IEEE Transactions on Knowledge and Data Engineering 8, 1 (1996), 157–172.
@@
- Kadavath et al. (2022) Saurav Kadavath, Tom Conerly, Amanda Askell, Tom Henighan, Dawn Drain, Ethan Perez, Nicholas Schiefer, Zac Hatfield-Dodds, Nova DasSarma, Eli Tran-Johnson, Scott Johnston, Sheer El-Showk, Andy Jones, Nelson Elhage, Tristan Hume, Anna Chen, Yuntao Bai, Sam Bowman, Stanislav Fort, Deep Ganguli, Danny Hernandez, Josh Jacobson, Jackson Kernion, Shauna Kravec, Liane Lovitt, Kamal Ndousse, Catherine Olsson, Sam Ringer, Dario Amodei, Tom Brown, Jack Clark, Nicholas Joseph, Ben Mann, Sam McCandlish, Chris Olah, and Jared Kaplan. 2022. Language Models (Mostly) Know What They Know. arXiv:2207.05221 [cs.CL] https://arxiv.org/abs/2207.05221
@@
- Kipf and Welling (2017) Thomas N. Kipf and Max Welling. 2017. Semi-Supervised Classification with Graph Convolutional Networks. In International Conference on Learning Representations. https://openreview.net/forum?id=SJU4ayYgl
- Koo et al. (2024) Terry Koo, Frederick Liu, and Luheng He. 2024. Automata-based constraints for language model decoding. In First Conference on Language Modeling. https://openreview.net/forum?id=BDBdblmyzY
- Kuchnik et al. (2023) Michael Kuchnik, Virginia Smith, and George Amvrosiadis. 2023. Validating Large Language Models with ReLM. In Proceedings of Machine Learning and Systems, D. Song, M. Carbin, and T. Chen (Eds.), Vol. 5. Curan, 457–476. https://proceedings.mlsys.org/paper_files/paper/2023/file/93c7d9da61ccb2a60ac047e92787c3ef-Paper-mlsys2023.pdf
@@
- Liu et al. (2025b) Chunwei Liu, Matthew Russo, Michael Cafarella, Lei Cao, Peter Baille Chen, Zui Chen, Michael Franklin, Tim Kraska, Samuel Madden, Rana Shahout, and Gerardo Vitagliano. 2025b. Palimpzest: Optimizing AI-Powered Analytics with Declarative Query Processing. In CIDR.
- Liu et al. (2025a) Hongfu Liu, Hengguan Huang, Xiangming Gu, Hao Wang, and Ye Wang. 2025a. On Calibration of LLM-based Guard Models for Reliable Content Moderation. In The Thirteenth International Conference on Learning Representations. https://openreview.net/forum?id=wUbum0nd9N
@@
- Madaan et al. (2023) Aman Madaan, Niket Tandon, Prakhar Gupta, Skyler Hallinan, Luyu Gao, Sarah Wiegreffe, Uri Alon, Nouha Dziri, Shrimai Prabhumoye, Yiming Yang, Shashank Gupta, Bodhisattwa Prasad Majumder, Katherine Hermann, Sean Welleck, Amir Yazdanbakhsh, and Peter Clark. 2023. SELF-REFINE: iterative refinement with self-feedback. In Proceedings of the 37th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’23). Curran Associates Inc., Red Hook, NY, USA, Article 2019, 61 pages.
- Malkov and Yashunin (2020) Yu A. Malkov and D. A. Yashunin. 2020. Efficient and Robust Approximate Nearest Neighbor Search Using Hierarchical Navigable Small World Graphs. IEEE Transactions on Pattern Analysis and Machine Intelligence 42, 4 (2020), 824–836. https://doi.org/10.1109/TPAMI.2018.2889473
- Marcus et al. (2011) Adam Marcus, Eugene Wu, David R. Karger, Samuel Madden, and Robert C. Miller. 2011. Crowdsourced Databases: Query Processing with People. In CIDR.
- OpenAI (2022) OpenAI. 2022. Introducing text and code embeddings. https://openai.com/index/introducing-text-and-code-embeddings/
- OpenAI (2024) OpenAI. 2024. Structured Outputs. https://platform.openai.com/docs/guides/structured-outputs/
@@
- Park et al. (2012) Hyunjung Park, Hector Garcia-Molina, Richard Pang, Neoklis Polyzotis, Aditya Parameswaran, and Jennifer Widom. 2012. Deco: a system for declarative crowdsourcing. Proc. VLDB Endow. 5, 12 (Aug. 2012), 1990–1993. https://doi.org/10.14778/2367502.2367555
@@
- Reimers and Gurevych (2019) Nils Reimers and Iryna Gurevych. 2019. Sentence-BERT: Sentence Embeddings using Siamese BERT-Networks. arXiv:1908.10084 [cs.CL] https://arxiv.org/abs/1908.10084
- Russo et al. (2025) Matthew Russo, Sivaprasad Sudhir, Gerardo Vitagliano, Chunwei Liu, Tim Kraska, Samuel Madden, and Michael Cafarella. 2025. Abacus: A Cost-Based Optimizer for Semantic Operator Systems. arXiv:2505.14661 [cs.DB] https://arxiv.org/abs/2505.14661
- Santhanam et al. (2024) Keshav Santhanam, Deepti Raghavan, Muhammad Shahir Rahman, Thejas Venkatesh, Neha Kunjal, Pratiksha Thaker, Philip Levis, and Matei Zaharia. 2024. ALTO: An Efficient Network Orchestrator for Compound AI Systems. In Proceedings of the 4th Workshop on Machine Learning and Systems (Athens, Greece) (EuroMLSys ’24). Association for Computing Machinery, New York, NY, USA, 117–125. https://doi.org/10.1145/3642970.3655844
@@
- TruLens (2021) TruLens. 2021. TruLens. https://www.trulens.org/
- Urban and Binnig (2024) Matthias Urban and Carsten Binnig. 2024. CAESURA: Language Models as Multi-Model Query Planners. In CIDR.
- Veličković et al. (2018) Petar Veličković, Guillem Cucurull, Arantxa Casanova, Adriana Romero, Pietro Liò, and Yoshua Bengio. 2018. Graph Attention Networks. In International Conference on Learning Representations. https://openreview.net/forum?id=rJXMpikCZ
- Wang and Li (2025) Jiayi Wang and Guoliang Li. 2025. AOP: Automated and Interactive LLM Pipeline Orchestration for Answering Complex Queries. In CIDR.
- Wehrstein et al. (2025) Johannes Wehrstein, Carsten Binnig, Fatma Özcan, Shobha Vasudevan, Yu Gan, and Yawen Wang. 2025. Towards Foundation Database Models. In CIDR.
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Brian Ichter, Fei Xia, Ed H. Chi, Quoc V. Le, and Denny Zhou. 2022. Chain-of-thought prompting elicits reasoning in large language models. In Proceedings of the 36th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’22). Curran Associates Inc., Red Hook, NY, USA, Article 1800, 14 pages.
- Widom (1996) Jennifer Widom. 1996. The starburst active database rule system. IEEE Transactions on Knowledge and Data Engineering 8, 4 (1996), 583–595.
@@
- Xu et al. (2019) Keyulu Xu, Weihua Hu, Jure Leskovec, and Stefanie Jegelka. 2019. How Powerful are Graph Neural Networks?. In International Conference on Learning Representations. https://openreview.net/forum?id=ryGs6iA5Km
- Zheng et al. (2023) Lianmin Zheng, Wei-Lin Chiang, Ying Sheng, Siyuan Zhuang, Zhanghao Wu, Yonghao Zhuang, Zi Lin, Zhuohan Li, Dacheng Li, Eric P. Xing, Hao Zhang, Joseph E. Gonzalez, and Ion Stoica. 2023. Judging LLM-as-a-judge with MT-bench and Chatbot Arena. In Proceedings of the 37th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’23). Curran Associates Inc., Red Hook, NY, USA, Article 2020, 29 pages.